OpenStudy (anonymous):

Factor completely: x2 − 4x + 9 is this prime

OpenStudy (anonymous):

Is what prime

OpenStudy (anonymous):

The answer

OpenStudy (anonymous):

No rearranging the polynomial we get, $$(x-2)^2=-5$$

OpenStudy (anonymous):

So its not even real

OpenStudy (anonymous):

(x − 1)(x − 9) (x − 3)(x − 3) (x + 3)(x − 3) Prime

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

$$(x-2)^2+5=x^2-4x+9$$

OpenStudy (anonymous):

Set $$(x-2)^2+5=0$$

OpenStudy (anonymous):

$$(x-2)^2=-5$$

OpenStudy (anonymous):

$$x-2=\pm \sqrt{-5}$$

OpenStudy (anonymous):

$$x=2 \pm \sqrt{-5}$$

OpenStudy (anonymous):

??

OpenStudy (anonymous):

look up "difference of squares"